Company A is selling 20 year bonds with a $50,000 par value and an annual coupon rate of 6.25%. The price of Company B's bonds is currently $11,200. They offer a $10,000 par value and mature in 10 years. The coupon rate is 8% and coupon payments are made semi-annually. Further research suggests to you that both bonds, which are rated as A+, introduce the same risk to you as an investor. If you bought a Company B bond for $10,450 and sold it two years later for $11,200 what would be your holding period rate of return?

ANSWER

1) 7.18%
2) 22.49%
3) $1,600
4) .2101
5) None of the above
Answer #1

Holding period return is the total return received from holding the bond over a period of time.

In the above case,

Purchase Price of Bond = $10,450

Selling Price = $ 11,200

Par Value of the Bond = $10,000

Coupon Rate (Annual) = 8%

Coupon payments are made semi-annually.

Hence, semi-annual coupon rate = 8% / 2 = 4%

Semi-annual coupon amount = 4% * 10,000 = $400

The Bond is held for a period of 2 years, hence total number of times coupon received = (2*2) = 4 times

Total Coupon received = $400 * 4 =$1,600

Following is the calculation of Holding Period Return :-

A C D Purchase Price of Bond B Total coupon received Selling Price of Bond B 10,450 1,600 11,200 Total Return Earned Holding

Hmo Purchase Price of Bond B 10450 Total coupon received -400*4 Selling Price of Bond B 11200 Total Return Earned Holding Per

Hence, Holding Period Return is 22.49%

Option 2 is correct